A poster is to have an area of 630 cm2 with 2.5 cm margins at the bottom and sides and a 5 cm margin at the top. Find the exact dimensions (in cm) that will give the largest printed area.


width  = cm


height = cm



1
Expert's answer
2022-05-22T14:57:32-0400

S=630 cm2

S=a*b, b=630/a

S= (2.5 *2 +a)*(7.5+b)


S= (5+a)*(7.5 + 630/a)= (37.5+3150/a+7.5a+630)=7.5a+3150/a+667.5

we find the first derivative and equate it to zero

S'= 7.5-3150a2=0

we solve the resulting equation and find the roots of the equation (i.e. the value of the unknown at which the value of the function (derivative) is 0). This will be the answer, i.e. the length of the side at which the area will be the largest.

7.5a2=3150

a2=420

a= 20.5; -20.5


width  = 20.5 cm


height = 30.7 cm


20.5 * 30.7 = 629.35 cm2